two incandescent lights are chosen at random from 19 lights of which 4 are defective. Find the probability that (a) none is defective? (b) exactly one is defective?


1
Expert's answer
2022-04-26T00:08:57-0400

There are "C_{19}^2=\frac{18\cdot19}{2}=9\cdot19=171" ways to choose "2" lights from "19". a). There are "C_{15}^2=\frac{14\cdot15}{2}=7\cdot 15=105" ways to choose "2" lights from "15" lights without defects. Thus, the probability is given by "p_1=\frac{C_{15}^2}{C_{19}^2}=\frac{105}{171}\approx0.614." b). There are "C_{15}^1=15" ways to choose "1" light without defect and "C_4^1=4" ways to choose "1" light with defect. Thus, the probability is as follows: "p_2=\frac{C_{15}^1C_4^1}{C_{19}^2}=\frac{60}{171}\approx0.351".

The answers are as follows: a) "p_1\approx0.614", b) "p_2\approx0.351."